OpenStudy (anonymous):

What is the horizontal shift for the graph of f(x) = (x + 5)3 - 7? A. Left 5 B. Right 5 C. Left 7 D. Right 7 I got a is that right?

OpenStudy (camper4834):

you are correct

OpenStudy (camper4834):

y=M(x-a)+b or y=mx+b

OpenStudy (anonymous):

ty